Scout, the dog, has epilepsy. His current medication, phenobarbital, is not controlling his seizures very well. He weighs 25 kg and is receiving a 100 mg tablet twice daily (BID). The therapeutic range for phenobarbital is 15 – 45 ug/ml. Therapeutic drug monitoring shows that Scout’s plasma concentration is 18 ug/ml. If the apparent Vd is 0.7 L/kg, and you want to get him to 30 ug/ml, how much more phenobarbital do you need to give to Scout?
Answer #1

Weight of the dog is 25kg

The dose of phenobarbital he is prescribed is 100mg tablet twice a day .

Therapeutic dose of phenobarbital is 15-45 microgram

Scout's plasma concentration of drug is 18 microgram /ml

VD is 0.7L/kg = 0.7×25= 17.5 L

We want to get his plasma concentration to 30microgram /ml

Formula :-

Peak concentration = dose / Vd

30 microgram /ml = dose / 17.5 L

Converting 30microgram / ml to Liter = 30× 0.001 = 0.03 microgram / L

dose = 0.03 microgram /L × 17.5L

Dose = 0.525microgram has to be added to the initial dose